# On decoding extended Han-Zhang codes

### v1

@@ Abstract
Extended Han-Zhang codes are a class of linear codes where each code is either a non-generalized Reed-Solomon (non-GRS) maximum distance separable (MDS) code or a near MDS (NMDS) code. They have important applications in communication, cryptography, and storage systems. While many algebraic properties and explicit constructions of extended Han-Zhang codes have been well studied in the literature, their decoding has been unexplored. In this paper, we focus on their decoding problems in terms of $\ell$-error-correcting pairs ($\ell$-ECPs) and deep holes. On the one hand, we determine the existence and specific forms of their $\ell$-ECPs, and further present an explicit decoding algorithm for extended Han-Zhang codes based on these $\ell$-ECPs, which can correct up to $\ell$ errors in polynomial time, with $\ell$ about half of the minimum distance. On the other hand, we determine the covering radius of extended Han-Zhang codes and characterize two classes of their deep holes, which are closely related to the maximum-likelihood decoding method. By employing these deep holes, we also construct more non-GRS MDS codes with larger lengths and dimensions, and discuss the monomial equivalence between them and the well-known Roth-Lempel codes. Some concrete examples are also given to support these results.
Index Terms:
Decoding, error-correcting pair, deep hole, non-GRS MDS code, NMDS code
@@ I Introduction
I-A Linear codes where each code is either non-GRS MDS or NMDS and their decoding problems
@@
Generalized Reed-Solomon (GRS) codes are a special class of MDS codes that have garnered a lot of attention due to their highly efficient encoding and decoding algorithms. Moreover, $n,k,n-k+1]_{q}$ GRS codes satisfy $n\leq q$, and together with the MDS conjecture $n\leq q+2$ [21], they realize nearly all possible parameter sets of MDS codes, except for the case $n>q$. However, it is still very interesting and vital to construct MDS codes that are not monomially equivalent to GRS codes, referred to as non-GRS MDS codes. These non-GRS MDS codes are important in classifying MDS codes, and are known to exhibit significant potential in resisting cryptographic attacks such as the Sidelnikov–Shestakov and Wieschebrink attacks, unlike GRS codes [27, 2]. Consequently, it is of particular interest to construct new families of linear codes in which each code is either a non-GRS MDS code or an NMDS code.
@@
There has been a substantial body of work [32, 33, 48, 17, 19, 13, 20, 57, 14, 43, 18, 45, 46, 31, 7, 54, 23, 28, 49] focusing on the construction of new families of codes that include both MDS and NMDS codes. In particular, the (+)-twisted generalized Reed-Solomon (TGRS) codes in [19], the ($\ast$)-TGRS codes in [20], the (+)-extended TGRS ((+)-ETGRS) codes in [57], and the extended Han–Zhang codes in [31] coincide with examples of linear code families where each code is either a non-GRS MDS code or an NMDS code.
@@
Of the codes in the previous paragraph, except for the extended Han–Zhang codes, decoding algorithms have been studied for the other three families of linear codes. Specifically, Beelen etal.et~al. [3] initially proposed a brute-force decoding method based on the established decoding algorithm of GRS codes when introducing TGRS codes. A simpler version was later presented in [4]. Although these decoding approaches are theoretically applicable for decoding (+)-TGRS and ($\ast$)-TGRS codes, they are not practical in real-world scenarios. As stated in [3] and [4], the former suffers from unexpected decoding complexity, while the latter lacks a rigorous proof ensuring that decoding works for any error vector up to the maximal decoding radius. Alternatively, He etal.et~al. [16] explored the existence of error-correcting pairs (ECPs) for (+)-TGRS and ($\ast$)-TGRS codes, and developed a corresponding decoding algorithm that operates in polynomial time. It is well-known that deep holes of linear codes are closely related to their maximum-likelihood decoding algorithms. Leveraging this insight, Fang etal.et~al. further identified the standard deep holes of (+)-TGRS codes of arbitrary lengths and explored the problem of determining all deep holes for full-length (+)-TGRS codes in [12]. Recently, Li etal.et~al. theoretically studied both the existence of ECPs and deep holes for (+)-ETGRS codes in [33]. Additionally, Jia etal.et~al. [22], Sui etal.et~al. [42], and Sun etal.et~al. [44] applied the Berlekamp–Massey algorithm for GRS codes and the extended Euclidean algorithm to propose efficient decoding algorithms for the dual codes of (+)-TGRS and ($\ast$)-TGRS codes. It is also worth noting from [19] and [20] that the dual codes of (+)-TGRS and ($\ast$)-TGRS codes may not be (+)-TGRS or ($\ast$)-TGRS codes again.
I-B Our motivations for decoding extended Han-Zhang codes based on ECPs and deep holes
@@
  • Fact 2: For the aforementioned Fact 1 codes, decoding algorithms based on ECPs and deep holes appear to be more practical and versatile than other known approaches.
  • Fact 3: Although the algebraic properties of extended Han–Zhang codes have been extensively studied in [31], their decoding problems remain unexplored. In particular, it was also proved in [31] that an extended Han-Zhang code is generally not monomially equivalent to an (+)-ETGRS code, due to different weight distributions.
The ECP decoding algorithm was introduced by Kötter [26] and Pellikaan [36] in 1992 (see Definition 5). It has also been demonstrated that ECPs exist for a wide range of linear codes, including cyclic codes, GRS codes, TGRS codes with single or double twists, lengthened GRS codes, Goppa codes, alternant codes, and algebraic geometry codes [10, 26, 36, 37, 15, 16, 17, 50, 33]. In particular, if a linear code possesses an $\ell$-ECP, this framework enables a universal algebraic decoding capable of correcting up to $\ell$ errors in polynomial time.
For another decoding algorithm related to deep holes (see Definition 6), it indeed represents an extreme case of the important maximum-likelihood decoding method and is consequently both interesting and difficult. By analyzing the structure of deep hole cosets, parameters of extended codes, connections with problems and results in finite geometry, syndromes, subset sum problems, full automorphism groups, and character sums, researchers explored deep holes of first-order Reed-Muller codes, GRS codes, (+)-TGRS codes, elliptic curve codes, and dual codes of Roth-Lempel codes in [35, 53, 9, 24, 25, 29, 58, 12, 33, 48] and the references therein. Furthermore, Pellikaan etal.et~al. explored the application of ECPs in the McEliece public-key cryptosystem for quantum-resistant cryptography [38], while Yang etal.et~al. [52] and Li etal.et~al. [33] respectively leveraged deep holes of binary BCH codes and ETGRS codes to derive new binary optimal codes and non-GRS MDS codes with larger lengths and dimensions. In consideration of the above facts and applications, an important and intriguing problem naturally arises:
@@
How can we decode extended Han-Zhang codes, particularly by leveraging their ECPs and deep holes? More specifically, is it possible to develop an efficient decoding algorithm for extended Han-Zhang codes based on their ECPs, and can we identify some of their deep holes?
@@ I-C Our contributions
    We use ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ to denote an $n+1,k]_{q}$ extended Han-Zhang code (see Definition 4 for more details). Let $\ell=\left\lfloor\frac{d({\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty))-1}{2}\right\rfloor$, where $d({\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty))$ denotes the minimum distance of the extended Han-Zhang code ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$. Utilizing structural properties, we identify extended Han-Zhang codes as subcodes of certain GRS codes and EGRS codes in Lemma 12. Building on this, we prove in Theorem 13.1) that ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ has only $(\ell-1)$-ECPs if it is an $n+1,k,n-k+2]_{q}$ MDS code and $2\nmid(n-k)$. We also show that ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ has $\ell$-ECPs in other cases by giving explicit forms of these ECPs directly in Theorems 13.2) and 14. More details on the exact values of $\ell$ and explicit forms of ECPs are summarized in Table I. Furthermore, we provide a decoding algorithm for ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ based on their ECPs in Theorem 15 and Algorithm 1, supported by specific decoding examples shown in Examples 1 and 2.
@@
    We prove that the covering radius of any extended Han-Zhang code is nk+1n-k+1 in Theorem 16 and establish a necessary and sufficient condition for a received vector to be a deep hole of ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ in Theorem 17. Leveraging these results, we further characterize when two classes of received vectors become deep holes of ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ in Theorems 19 and 20, where the first class is associated with the polynomial $g(x)=g_{k-1}x^{k-1}+f(x)$ satisfying $g_{k-1}\in\mathbb{F}_{q}^{*}$ and $f(x)\in\mathcal{V}_{k}$, the second one is related to the polynomial $g(x)=g_{k+1}x^{k+1}+g_{k-1}x^{k-1}+f(x)$ with $g_{k+1}\in\mathbb{F}_{q}^{*}$, $g_{k-1}\in\mathbb{F}_{q}$, and $f(x)\in\mathcal{V}_{k}$, and $\mathcal{V}_{k}=\{f(x)=\sum_{i=0}^{k-2}f_{i}x^{i}+f_{k}x^{k}:~f_{i}\in\mathbb{F}_{q},~\forall~i\in\{0,\ldots,k-2,k\}\}$.
@@
    Applying the deep holes obtained and the non-GRS MDS properties of extended Han-Zhang codes, we construct more $n+2,k+1,n-k+2]_{q}$ non-GRS MDS codes and analyze their monomial equivalence with the well-known Roth-Lempel codes in Theorems 21 and 22 as well as Remark 3. Notably, our results also confirm that there is a special family of Roth-Lempel codes where each code is either non-GRS MDS or NMDS, as stated in Remark 4. Additionally, we provide an efficient algorithm to generate these non-GRS MDS codes with larger lengths and dimensions in Theorem 23 and Algorithm 2, and validate our findings with specific examples in Examples 3 and 4.
This paper is organized as follows. After the introduction, Section II reviews fundamental notation and results on GRS codes, Roth-Lempel codes, extended Han-Zhang codes, ECPs, and deep holes. Section III establishes the existence and explicit forms of $\ell$-ECPs for extended Han-Zhang codes, and presents a decoding algorithm based on their $\ell$-ECPs. Section IV explores the deep holes of extended Han-Zhang codes and leverages these findings to construct more non-GRS MDS codes with larger lengths and dimensions. We also discuss the monomial equivalence between these newly derived non-GRS MDS codes and Roth-Lempel codes. Finally, Section V concludes this paper.
@@ II Preliminaries
In this section, we recall some basic notation and results on GRS codes, Roth-Lempel codes, extended Han-Zhang codes, ECPs, and deep holes.
@@ II-A Basic notation
  • For any vectors ${\bf x}=(x_{1},\ldots,x_{n})\in\mathbb{F}_{q}^{n}$ and ${\bf y}=(y_{1},\ldots,y_{n})\in\mathbb{F}_{q}^{n}$, their standard product and Schur product are respectively defined by
@@
    $\mathcal{C}\star\mathcal{D}=\{(c_{1}d_{1},\ldots,c_{n}d_{n}):~(c_{1},\ldots,c_{n})\in\mathcal{C}~{\rm and}~(d_{1},\ldots,d_{n})\in\mathcal{D}\}.$
@@
II-B GRS codes, Roth-Lempel codes, and extended Han-Zhang codes
@@
$G_{{\mathrm{GRS}}_{k}({\mathcal{S}},\mathbf{v})}=\begin{pmatrix}v_{1}&\ldots&v_{n}\\ v_{1}a_{1}&\ldots&v_{n}a_{n}\\ \vdots&\ddots&\vdots\\ v_{1}a_{1}^{k-1}&\ldots&v_{n}a_{n}^{k-1}\\ \end{pmatrix}~{\rm and}~G_{{\mathrm{GRS}}_{k}({\mathcal{S}},\mathbf{v},\infty)}=\begin{pmatrix}v_{1}&\ldots&v_{n}&0\\ v_{1}a_{1}&\ldots&v_{n}a_{n}&0\\ \vdots&\ddots&\vdots&\vdots\\ v_{1}a_{1}^{k-1}&\ldots&v_{n}a_{n}^{k-1}&1\\ \end{pmatrix}$, (1)
@@
( [11, Lemma 5] and [21, Theorem 5.3.3]) The following statements hold.
@@
    ${\mathrm{GRS}}_{k}({\mathcal{S}},\mathbf{v},\infty)^{\perp}={\mathrm{GRS}}_{n-k+1}({\mathcal{S}},\mathbf{v}^{-1}\star{{\mathbf{u}}},\infty){\mathrm{diag}}(\underbrace{1,\ldots,1}_{n},-1)$.
@@
( [39]) Let $k$ and $n$ be two positive integers satisfying $3\leq k\leq n\leq q$. For any $\delta\in\mathbb{F}_{q}$, a Roth-Lempel code, denoted by ${\mathrm{RL}}_{k,\delta}({\mathcal{S}})$, is an $n+2,k]_{q}$ linear code defined by
$\begin{split}{\mathrm{RL}}_{k,\delta}({\mathcal{S}})=\{(f(a_{1}),\ldots,f(a_{n}),f_{k-1},f_{k-2}+\delta f_{k-1}):~f(x)\in\mathbb{F}_{q}[x]_{k}\},\end{split}$ (2)
where $f_{k-1}$ and $f_{k}$ are the coefficients of $x^{k-1}$ and $x^{k}$ in $f(x)$, respectively. Moreover, it has a generator matrix of the form
$G_{{\mathrm{RL}}_{k,\delta}({\mathcal{S}})}=\left(\begin{array}[]{cccccc}1&\ldots&1&0&0\\ a_{1}&\ldots&a_{n}&0&0\\ \vdots&\ddots&\vdots&\vdots&\vdots\\ a^{k-3}_{1}&\ldots&a^{k-3}_{n}&0&0\\ a^{k-2}_{1}&\ldots&a^{k-2}_{n}&0&1\\ a^{k-1}_{1}&\ldots&a^{k-1}_{n}&1&\delta\\ \end{array}\right).$ (9)
Definition 3.
For a subset ${\mathcal{S}}\subseteq\mathbb{F}_{q}$, we call $\mathcal{S}$} an $(n,k,\delta)$-set if
$|{\mathcal{S}}|=n~{\rm and}~a_{i_{1}}+\ldots+a_{i_{k}}\neq\delta~{\rm for~any}~\{a_{i_{1}},\ldots,a_{i_{k}}\}\subseteq{\mathcal{S}}.$
Moreover, $\mathcal{S}$} is an $(n,k,0)$-set if and only if $\mathcal{S}$} is $k$-zero-sum free. Conversely, we say that $\mathcal{S}$} contains a $k$-zero-sum subset if $\mathcal{S}$} is not $k$-zero-sum free.
In [39], Roth and Lempel established a necessary and sufficient condition for a Roth-Lempel code ${\mathrm{RL}}_{k,\delta}({\mathcal{S}})$ to be non-GRS MDS as follows.
Lemma 2.
( [39, Page 655]) For any $3\leq k\leq n\leq q$ and $\delta\in\mathbb{F}_{q}$, a Roth-Lempel code ${\mathrm{RL}}_{k,\delta}({\mathcal{S}})$ is an $n+2,k,n-k+3]_{q}$ non-GRS MDS code if and only if $\mathcal{S}$} is an $(n,k-1,\delta)$-set.
Definition 4.
( [31, Definition 1]) Let $n$ and $k$ be two positive integers satisfying $3\leq k\leq n-2\leq q-2$. An extended Han-Zhang code, denoted by ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ is an $n+1,k]_{q}$ linear code
@@
The following lemmas recall some basic properties of extended Han-Zhang codes obtained in [31].
@@
( [31, Theorem \@slowromancapiii@.2]) The extended Han-Zhang code ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ has a generator matrix and a parity-check matrix of the form
$G_{{\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)}=\begin{pmatrix}v_{1}&\ldots&v_{n}&0\\ v_{1}a_{1}&\ldots&v_{n}a_{n}&0\\ \vdots&\ddots&\vdots&\vdots\\ v_{1}a_{1}^{k-2}&\ldots&v_{n}a_{n}^{k-2}&0\\ v_{1}a_{1}^{k}&\ldots&v_{n}a_{n}^{k}&1\\ \end{pmatrix}~{\rm and}~H_{{\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)}=\begin{pmatrix}\frac{u_{1}}{v_{1}}&\ldots&\frac{u_{n}}{v_{n}}&0\\ \frac{u_{1}}{v_{1}}a_{1}&\ldots&\frac{u_{n}}{v_{n}}a_{n}&0\\ \vdots&\ddots&\vdots&\vdots\\ \frac{u_{1}}{v_{1}}a_{1}^{n-k-2}&\ldots&\frac{u_{n}}{v_{n}}a_{n}^{n-k-2}&0\\ \frac{u_{1}}{v_{1}}a_{1}^{n-k-1}&\ldots&\frac{u_{n}}{v_{n}}a_{n}^{n-k-1}&-1\\ \frac{u_{1}}{v_{1}}a_{1}^{n-k}&\ldots&\frac{u_{n}}{v_{n}}a_{n}^{n-k}&-\sum_{i=1}^{n}a_{i}\end{pmatrix}$, (10)
@@
( [31, Theorem \@slowromancapiii@.4]) Any extended Han-Zhang code ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ is either MDS or NMDS. Furthermore, the following statements hold.
@@
( [31, Theorem \@slowromancapiii@.7]) Any extended Han-Zhang code ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ is non-GRS.
II-C ECPs and deep holes
This subsection reviews some basic definitions of ECPs and deep holes. For any linear code $\mathcal{C}$}, a pair of linear codes $(\mathcal{A},\mathcal{B})$ is an $\ell$-error-correcting pair ($\ell$-ECP) of $\mathcal{C}$} that can yield a decoding algorithm capable of uniquely correcting up to $\ell$ errors in polynomial time. From [36, Corollary 2.15], we have $\ell\leq\lfloor\frac{d({\mathcal{C}})-1}{2}\rfloor$. The following is an alternative definition of the $\ell$-ECP.

### v2

# Unique Decoding of Extended Subcodes of GRS Codes Using Error-Correcting Pairs

@@ Abstract
Extended subcodes of generalized Reed-Solomon (ESGRS) codes are a class of linear codes where each code is either a non-GRS maximum distance separable (MDS) code or a near MDS (NMDS) code. They have important applications in distributed storage systems and cryptography. While many algebraic properties and explicit constructions of ESGRS codes have been well studied in the literature, their decoding has been unexplored. In this paper, we focus on their unique decoding problems in terms of $\ell$-error-correcting pairs ($\ell$-ECPs). We determine the existence and specific forms of their $\ell$-ECPs by utilizing GRS codes and their variants. We further present an explicit decoding algorithm for ESGRS codes based on these $\ell$-ECPs, which can uniquely correct up to $\ell$ errors in polynomial time, with $\ell$ about half of the minimum distance. Compared with other potential decoding approaches, our algorithm is more efficient in certain scenarios. Some concrete examples are also given to support these results.
@@ I Introduction
I-A Linear codes where each code is either non-GRS MDS or NMDS and decoding problems
@@
Generalized Reed-Solomon (GRS) codes are a special class of MDS codes that have garnered a lot of attention due to their highly efficient encoding and decoding algorithms. Moreover, $n,k,n-k+1]_{q}$ GRS codes satisfy $n\leq q$, and together with the MDS conjecture $n\leq q+1$ in general except for two special cases [17], they realize nearly all possible parameter sets of MDS codes, except for the case $n>q$. However, it is still very interesting and vital to construct MDS codes that are not monomially equivalent to GRS codes, referred to as non-GRS MDS codes. These non-GRS MDS codes are important in classifying MDS codes, and are known to exhibit significant potential in resisting cryptographic attacks such as the Sidelnikov–Shestakov and Wieschebrink attacks, unlike GRS codes [21, 2]. Consequently, it is of particular interest to construct new families of linear codes in which each code is either a non-GRS MDS code or an NMDS code.
@@
There has been a substantial body of work [26, 42, 14, 15, 11, 16, 47, 12, 36, 38, 39, 25, 6, 45, 19, 22, 43] focusing on the construction of new families of codes that include both MDS and NMDS codes. In particular, the (+)-twisted generalized Reed-Solomon ((+)-TGRS) codes in [15], the ($\ast$)-TGRS codes in [16], the (+)-extended TGRS ((+)-ETGRS) codes in [47], and the extended subcodes of GRS (ESGRS) codes in [25] coincide with examples of linear code families where each code is either a non-GRS MDS code or an NMDS code.
As Huffman and Pless emphasized in [17], finding efficient (fast) decoding algorithms is a major area of research in coding theory because of their practical applications. There are many different decoding approaches, such as unique decoding and non-unique decoding, which can be classified according to their decoding outputs. Compared with non-unique decoding, unique decoding guarantees both correctness and unambiguity of the decoding output within its designed decoding radius [10]. Specifically, when the number of errors does not exceed half of the minimum distance, unique decoding deterministically recovers the transmitted codeword and excludes any decoding ambiguity. This property is particularly crucial in cryptographic and security-sensitive applications, where multiple candidate outputs are unacceptable [27, 32]. Therefore, in this paper, we focus on the unique decoding of such families of linear codes where each code is either non-GRS MDS or NMDS. For convenience, unless stated otherwise, all subsequent references to “decoding” refer to unique decoding.
I-B Our motivations for unique decoding ESGRS codes based on ECPs
@@
Of the codes in the previous subsection, except for the ESGRS codes, unique decoding algorithms have been studied for the other three families of linear codes. Let $q$ be the alphabet size and $n$ be the code length with $n\leq q$. Specifically, Beelen etal.et~al. [3] initially proposed a brute-force decoding method based on the established decoding algorithm of GRS codes when introducing TGRS codes with time complexity $O(q^{t}n^{2})$ [37], where $t$ is the number of twists. A simpler version was later presented in [4]. Although these decoding approaches are theoretically applicable for decoding (+)-TGRS and ($\ast$)-TGRS codes, they are not practical in real-world scenarios. As stated in [3] and [4], the former suffers from unexpected decoding complexity, while the latter lacks a rigorous proof ensuring that decoding works for any error vector up to the maximal decoding radius.
Additionally, Jia etal.et~al. [18], Sui etal.et~al. [35], and Sun etal.et~al. [37] applied the Berlekamp–Massey algorithm for GRS codes and the extended Euclidean algorithm to propose efficient decoding algorithms for the dual codes of MDS (+)-TGRS and MDS ($\ast$)-TGRS codes. Wang etal.et~al. [41] further extended it to the NMDS cases. According to [37], this method has time complexity $O(qn)$ in general. It is also worth noting from [15] and [16] that the dual codes of (+)-TGRS and ($\ast$)-TGRS codes may not be (+)-TGRS or ($\ast$)-TGRS codes again. Consequently, this method cannot be straightforwardly applied to arbitrary (+)-TGRS and ($\ast$)-TGRS codes.
Alternatively, He etal.et~al. [13] explored the forms of $\ell$-error-correcting pairs ($\ell$-ECPs) for general (+)-TGRS and ($\ast$)-TGRS codes. Li etal.et~al. theoretically studied the existence of $\ell$-ECPs for MDS (+)-ETGRS codes in [26] and Li etal.et~al. provided explicit constructions of $\ell$-ECPs for NMDS (+)-ETGRS codes in [24], where $\ell$ is approximately half of the minimum distance. It can be observed that the resulting decoding algorithms can uniquely correct up to $\ell$ errors in polynomial time on $n$, independent of the alphabet size $q$, thereby making them much faster than other known approaches in certain scenarios.
@@
  • Fact 2: For the aforementioned Fact 1 codes, decoding algorithms based on ECPs appear to be more practical and versatile than other known approaches.
  • Fact 3: Although the algebraic properties of ESGRS codes have been extensively studied in [25], their decoding problems remain unexplored. In particular, it was also proved in [25] that an ESGRS code is generally not monomially equivalent to a TGRS code or its extended code, due to different weight distributions.
The concept of ECP was introduced by Kötter [20] and Pellikaan [29] in 1992 (see Definition 3). It has also been further demonstrated that ECPs exist for a wide range of linear codes, including cyclic codes, GRS codes, TGRS codes with single or double twists, lengthened GRS codes, Goppa codes, alternant codes, and algebraic geometry codes [8, 20, 29, 30, 13, 14, 44, 26]. In particular, if a linear code possesses an $\ell$-ECP, this framework enables a universal algebraic decoding capable of uniquely correcting up to $\ell$ errors in polynomial time, where the specific time complexity depends on the related extension degree of finite fields [30, 8]. Furthermore, Pellikaan etal.et~al. explored the application of ECPs in the McEliece public-key cryptosystem for quantum-resistant cryptography [31]. In consideration of the above facts and applications, an important and intriguing problem naturally arises:
@@
How can we decode ESGRS codes by leveraging their ECPs? More specifically, is it possible to develop an efficient decoding algorithm for ESGRS codes based on their ECPs?
@@ I-C Our contributions
    We use ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ to denote an $n+1,k]_{q}$ ESGRS code (see Definition 2 for more details). Let $\ell=\left\lfloor\frac{d({\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty))-1}{2}\right\rfloor$, where $d({\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty))$ denotes the minimum distance of the ESGRS code ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$. Utilizing structural properties, we identify ESGRS codes as subcodes of certain GRS codes and EGRS codes in Lemma 8. Building on this, we prove in Theorem 9.1) that ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ has only $(\ell-1)$-ECPs if it is an $n+1,k,n-k+2]_{q}$ MDS code and $2\nmid(n-k)$. In this case, the ECP can be determined by two GRS codes.
@@
    We also show that ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ has $\ell$-ECPs in other cases by providing explicit constructions of these ECPs in Theorems 9.2) and 10, based on variants of GRS codes. More details on the exact values of $\ell$ and explicit forms of ECPs are summarized in Table I.
@@
    We provide an explicit decoding algorithm for ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ with time complexity $O(n^{3})$ based on their ECPs in Theorem 11 and Algorithm 1. We further compare our decoding algorithm with other potential approaches and identify some explicit advantages in Subsection IV-B. We also present two specific decoding examples shown in Examples 1 and 2 to illustrate our results.
This paper is organized as follows. After the introduction, Section II reviews fundamental notation and results on GRS codes, ESGRS codes, and ECPs. Section III establishes the existence and explicit forms of $\ell$-ECPs for ESGRS codes. Section IV presents an explicit decoding algorithm based on their $\ell$-ECPs and further discusses its complexity. Finally, Section V concludes this paper.
@@ II Preliminaries
In this section, we recall some basic notation and results on GRS codes, ESGRS codes, and ECPs.
@@ II-A Basic notation
  • For any vectors ${\bf x}=(x_{1},\ldots,x_{n})\in\mathbb{F}_{q}^{n}$ and ${\bf y}=(y_{1},\ldots,y_{n})\in\mathbb{F}_{q}^{n}$, their ordinary inner product and Schur product are respectively defined by
@@
    $\mathcal{C}\star\mathcal{D}=\{(c_{1}d_{1},$ $\ldots,c_{n}d_{n}):$
    $(c_{1},\ldots,c_{n})\in\mathcal{C},~(d_{1},\ldots,d_{n})\in\mathcal{D}\}.$
@@
II-B GRS codes and ESGRS codes
@@
$G_{{\mathrm{GRS}}_{k}({\mathcal{S}},\mathbf{v})}=\begin{pmatrix}v_{1}&\ldots&v_{n}\\ v_{1}a_{1}&\ldots&v_{n}a_{n}\\ \vdots&\ddots&\vdots\\ v_{1}a_{1}^{k-1}&\ldots&v_{n}a_{n}^{k-1}\\ \end{pmatrix}$ (1)
and
$G_{{\mathrm{GRS}}_{k}({\mathcal{S}},\mathbf{v},\infty)}=\begin{pmatrix}v_{1}&\ldots&v_{n}&0\\ v_{1}a_{1}&\ldots&v_{n}a_{n}&0\\ \vdots&\ddots&\vdots&\vdots\\ v_{1}a_{1}^{k-1}&\ldots&v_{n}a_{n}^{k-1}&1\\ \end{pmatrix}$, (2)
@@
( [9, Lemma 5] and [17, Theorem 5.3.3]) The following statements hold.
@@
    ${\mathrm{GRS}}_{k}({\mathcal{S}},\mathbf{v},\infty)^{\perp}={\mathrm{GRS}}_{n-k+1}({\mathcal{S}},\mathbf{v}^{-1}\star{{\mathbf{u}}},\infty)\cdot A$ with $A={\mathrm{diag}}(\underbrace{1,\ldots,1}_{n},-1)$.
@@
( [25, Definition 1]) Let $n$ and $k$ be two positive integers satisfying $3\leq k\leq n-2\leq q-2$. An extended subcode of GRS (ESGRS) code, denoted by ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ is an $n+1,k]_{q}$ linear code
@@
The following lemmas recall some basic properties of ESGRS codes obtained in [25].
@@
( [25, Theorem III.2]) The ESGRS code ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ has a generator matrix and a parity-check matrix of the form
$G_{{\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)}=\begin{pmatrix}v_{1}&\ldots&v_{n}&0\\ v_{1}a_{1}&\ldots&v_{n}a_{n}&0\\ \vdots&\ddots&\vdots&\vdots\\ v_{1}a_{1}^{k-2}&\ldots&v_{n}a_{n}^{k-2}&0\\ v_{1}a_{1}^{k}&\ldots&v_{n}a_{n}^{k}&1\\ \end{pmatrix}$ (3)
and
$H_{{\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)}=\begin{pmatrix}\frac{u_{1}}{v_{1}}&\ldots&\frac{u_{n}}{v_{n}}&0\\ \frac{u_{1}}{v_{1}}a_{1}&\ldots&\frac{u_{n}}{v_{n}}a_{n}&0\\ \vdots&\ddots&\vdots&\vdots\\ \frac{u_{1}}{v_{1}}a_{1}^{n-k-2}&\ldots&\frac{u_{n}}{v_{n}}a_{n}^{n-k-2}&0\\ \frac{u_{1}}{v_{1}}a_{1}^{n-k-1}&\ldots&\frac{u_{n}}{v_{n}}a_{n}^{n-k-1}&-1\\ \frac{u_{1}}{v_{1}}a_{1}^{n-k}&\ldots&\frac{u_{n}}{v_{n}}a_{n}^{n-k}&-\sum_{i=1}^{n}a_{i}\end{pmatrix}$, (4)
@@
For a subset ${\mathcal{S}}\subseteq\mathbb{F}_{q}$, we say that $\mathcal{S}$} contains a $k$-zero-sum subset if there exists a nonempty subset $\{a_{i_{1}},\ldots,a_{i_{k}}\}\subseteq{\mathcal{S}}$ such that $a_{i_{1}}+\ldots+a_{i_{k}}=0$. Conversely, we say that $\mathcal{S}$} is $k$-zero-sum free.
@@
( [25, Theorem III.4]) Any ESGRS code ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ is either MDS or NMDS. Furthermore, the following statements hold.
@@
( [25, Theorem III.7]) Any ESGRS code ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ is non-GRS.
II-C ECPs of linear codes
This subsection reviews some basic definitions of ECPs . For any linear code $\mathcal{C}$}, a pair of linear codes $(\mathcal{A},\mathcal{B})$ is an $\ell$-error-correcting pair ($\ell$-ECP) of $\mathcal{C}$} that can yield a decoding algorithm capable of uniquely correcting up to $\ell$ errors in polynomial time. From [29, Corollary 2.15], we have $\ell\leq\lfloor\frac{d({\mathcal{C}})-1}{2}\rfloor$. The following is an alternative definition of the $\ell$-ECP.